6. The advertising elasticity of a firm is 1.5 as advertising expenditure increases from $10 to $12 million. If demand is 50 at an advertising expenditure of $12 million, what will be the demand at an advertising expenditure of $10 million? 


1
Expert's answer
2022-02-24T11:39:01-0500

solution: Let p1=12. p2=10. Q1=50. we must to find Q2

"E=\frac{Q2-Q1}{P2-P1}\times\frac{P2+P1}{Q2+Q1}"

"1.5=\frac{50-Q1}{12-10}\times\frac{12+10}{50+Q1}" ="15\times(15+Q1)=11\times(50-Q1)"

="75+15Q1=550-11Q1"

12.5Q1=475

Q1=38